(33%) Problem 3: A uniformly dense 3.5 kg rod with a length L :: 2 m is located in the presence of the following gravitational field: kgm a> what is the magnitude of the gravitational force (in nN) acting on the rod?
Answer #1

Elemental max = 3.5/2 dx = 1.75 dx

dF = (1.75 dx) (3 x 10^-9)x^2

=(5. 25 x 10^-9)x^2 dx

F = lim(0->2)(5.25 x 10^-9)(x^3/3)

F = (5.25 x 10^-9)(8/3)

F = 1.4 x 10^-8 = 14 nN